What reaction do I use for this (there aren't any given)
In a softdrink the proton ion, H concentration was measured to be [H 9.83 x 10 M. What is the hydroxide ion concentration? [oH
Answer #1

[H+] = 9.83 x 10-3 M

pH = -log[H+] = -log[ 9.83 x 10-3 ] = 2

pH= 2

pH + pOH = 14

pOH =14-2 = 12

[OH ] = 10-12 M